OpenStudy (love2laugh81):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (bobo-i-bo):

have you learnt what the dot product is?

OpenStudy (dumbcow):

\[\cos \theta = \frac{u*v}{|u| |v|}\]